282006:

Steve set up a business making friendship bracelets. He earned $70 in January. He is projected to earn 120% more in February. How much is Steve projected to make in February?

NonportritV2:

Solve the question step by step to get your answer: 120% of 70 First find the answer to this: 70/x=100/120 (70/x)*x=(100/120)*x 70=0.833333333333*x (0.833333333333) to get x 70/0.833333333333=x 84=x x=84 $70+ $84 = 154$ In conclusion the best answer for this question is...... "154$." Steve would have a 120% increase of what he usually makes and if 120% of 70 is 84 then add 70 to 84 to get 154 which is your answer.